Paul’s Plumbing charges $55 to visit a job site plus $35 per hour for labor. Reliable Plumbing charges $75 to visit a job site plus $30 per hour for labor. For how many hours of labor would the total charge be the same for the two companies?
Answer by Theo(13342)

let x = number of hours.

55 + 35*x = 75 + 30*x

subtract 30*x from both sides of the equation and subtract 55 from both sides of the equation to get:

5*x = 20

divide both sides of the equation by 5 to get:

x = 4

the total charge should be the same in 4 hours.

55 + 35*4 = 55 + 140 = 195

75 + 30*4 = 75 + 120 = 195

4 hours is confirmed.

that's your answer.

total charge will be the same for both companies when the number of hours for the job is equal to 4.